OpenStudy (anonymous):

Of all the soft drink consumers in a particular sales region, 30% prefer Brand A and 70% prefer Brand B. Of all these soft drink consumers, 20% prefer Brand A and are female, and 40% prefer Brand B and are female. What is the probability that a randomly selected consumer is female, given that the person prefers Brand A? 0.18 0.21 0.34 0.67

OpenStudy (anonymous):

@beginnersmind

OpenStudy (anonymous):

@karatechopper

OpenStudy (anonymous):

are u sure?

OpenStudy (anonymous):

ok thanks though

OpenStudy (anonymous):

hey umm can i ask u something?

OpenStudy (anonymous):

did u copy and paste what you just replied to me?

OpenStudy (anonymous):

this : 0.3 * 0.2= 0.06. Then you do 0.7 * 0.4= 0.28. Finally, 0.06/0.28= 0.21

OpenStudy (anonymous):

lol yeah what i actually got was .67 but when i looked it up it showed that

OpenStudy (anonymous):

do u think it might be 67?

OpenStudy (anonymous):

i looked it up too and theres a comment that says that that's not the right answer

OpenStudy (anonymous):

hmm i did it my way and got .67

OpenStudy (anonymous):

omg im scared lol i dont even know what to put!

OpenStudy (anonymous):

are you taking Flvs?

OpenStudy (anonymous):

haha idk what that is lol im doing plato to do credit recovery

OpenStudy (anonymous):

i failed geometry and now im stuck doing this

OpenStudy (anonymous):

well email your online teacher then ask him/her for help.

OpenStudy (anonymous):

be honest, do u think it's 67?

OpenStudy (anonymous):

yes but if someone says its not them maybe im wrong.

OpenStudy (anonymous):

no, the comment said that 0.21 is wrong

OpenStudy (anonymous):

oh well then i guess i might be right. only one way to find out

OpenStudy (anonymous):

ok im gonna choose 67

OpenStudy (anonymous):

u were right!!!

OpenStudy (anonymous):

YAY lol

OpenStudy (anonymous):

thank you so much! :D

OpenStudy (anonymous):

your welcome :)